How many liters of a 50% acid solution must be mixed with a 10% acid solution to get 360 L of a 40% acid solution?
Answer by mananth(16946)

Acid percent ---------------- quantity
Type I 50.00% ---------------- x liters
Type II 10.00% ------ 360 - x liters
Mixture 40.00% ---------------- 360

50.00%x +10.00%(360-x)= 40.00% *360
50x+10(360-x)=14400
50x+3600-10x=14400
50x-10x =14400-3600
40x= 10800
/ 40
x=270 liters 50.00%
90 liters 10.00%
m.ananth@hotmail.ca